OpenStudy (anonymous):

x^2 + 14x-4=0 Solve.

OpenStudy (anonymous):

That left side does not look factorable so you can use the Quadratic Formula: \[\large x= \frac{-b \pm \sqrt{b^2-(4)(a)(c)}}{2(a)}\] where: a=1 b=14 c=-4

OpenStudy (anonymous):

thanks